Passing through point (3,7)and parallel to the line with equation 5x+3y=12
Answer by ewatrrr(24785)

 
Hi,


Passing through point (3,7)and parallel to the line with equation 5x+3y=12
parallel to the line with equation
5x+3y=12 OR y = (-5/3)x + 4 m = (-5/3)
Parallel lines have identical slopes
New Line
y = (-5/3)x + b |Using ordered pair (3,7) to solve for b
7 = (-5/3)*3 + b
12 = b
y = (-5/3)x + 12 OR 5x +3y = 36